### v1

# A Polyharmonic Liouville Hierarchy on Complete Manifolds of Nonnegative Ricci Curvature

Abstract.
In this paper, we establish a complete Liouville–type hierarchy for polyharmonic functions on a broad class of Riemannian manifolds with nonnegative Ricci curvature. Extending Yau’s classical result for harmonic functions and our recent biharmonic Liouville theorem, we prove that on any complete manifold of nonnegative Ricci curvature with a pole, every kk–polyharmonic function of growth $o(r^{2(k-1)})$ must in fact be (k1)(k-1)–polyharmonic. Iterating this procedure yields the result that all polyharmonic functions of sublinear growth are constant. The key innovation is a new $L^{2}$ estimate for the Laplacian of a polyharmonic function, obtained by induction through a delicate cutoff construction combined with a hole–filling argument. This provides the first sharp geometric extension of the Euclidean classification of polyharmonic functions to manifolds of nonnegative Ricci curvature, and completes a natural hierarchy of Yau–type Liouville theorems for iterates of the Laplacian.
Key words: Liouville theorem; polyharmonic functions; polynomial growth
AMS 2020 Mathematics subject classification: Primary 31C05, 53C21
1. Introduction
Harmonic functions have been a central object of study in Geometric Analysis. Among the most celebrated results, and a foundational one, is Yau’s Liouville theorem [18], from 1975, which asserts that on a complete Riemannian manifold of nonnegative Ricci curvature, any bounded harmonic function must be constant.
Since Yau, the number of interesting and important results about harmonic functions and spaces of harmonic functions have been constantly increasing. Among them, just to name a few, there are the beautiful works of Li and Tam [13], Colding and Minicozzi [7, 8] on the structure of the spaces of harmonic functions of polynomial growth, Ni and Tam [15] on pluriharmonic functions, and more recently, the work of Liu for holomorphic functions in [14].
However, what might seem a natural question has received little attention in the literature. Although the intrinsic interest the study of higher order elliptic operators may have (for instance, the Paneitz operator in conformal, or the study of polyharmonic flows for curves) there are very few results regarding their Liouville-type properties. As a few interesting exceptions we refer to the works of Branding [2, 3], where the author considers biharmonic maps, and proves Liouville-type theorems under the smallness of certain Sobolev norms. Recently the authors proved the following theorem (see [6]).
Theorem 1.
Let $(M,g)$ be a complete Riemannian manifold with a pole and such that $\mbox{Ric}_{g}\geq 0$. Then any biharmonic function $u\in C^{4}\left(M\right)$ of subquadratic growth is harmonic. Therefore, any biharmonic function of sublinear growth must be constant.
The purpose of this paper is to extend the previous result to $k$-polyharmonic functions. Recall that a $k$-polyharmonic function on a Riemannian manifold $\left(M,g\right)$ is a function $u:M\longrightarrow\mathbb{R}$ is a function $u\in C^{2k}\left(M\right)$ such that
$\Delta^{k}u=0,$
where $\Delta$ is the Laplacian.
In this paper, we shall consider manifolds with a pole, that is manifolds $M$ for which there is a point $p\in M$ such that the exponential map
$\mbox{exp}_{p}:\,T_{p}M\longrightarrow M$
is a diffeomorphism We are ready to state our main result.
Theorem 2.
Let $(M,g)$ be a complete Riemannian manifold with a pole and such that $\mbox{Ric}_{g}\geq 0$. Then any $k$-polyharmonic function $u\in C^{2k}\left(M\right)$ with
$u=o(r^{2(k-1)})\quad\mbox{as}\quad r\rightarrow\infty,$
where $r=d\left(p,x\right)$, is (k1)(k-1)-polyharmonic. Therefore, any polyharmonic function of sublinear growth must be constant.
As the reader may know, Theorem 2 has been known in the case of $\mathbb{R}^{n}$ with the Euclidean metric for quite some time: the critical growth separating genuinely $k$-polyharmonic behavior from lower-order behavior is $r^{2(k-1)}$. This shows that our theorem is sharp and is a natural extension of what is known in the Euclidean setting.
We now give an account of the subtleties in the proof of Theorem 2. First, we review the proof of Theorem 1. Our proof in [6] used an energy estimate for the $L^{2}$-norm of $\Delta u$ and a Cacciopoli-type inequality
(1) $\int_{B_{s}}\left|\nabla u\right|^{2}\,dV\leq\frac{C}{\left(r-s\right)^{2}}\int_{B_{r}}u^{2}\,dV$
valid for both harmonic and biharmonic functions. All this was used to prove
(2) $\int_{B_{s}}\left|\Delta u\right|^{2}\,dV\leq\frac{C}{\left(r-s\right)^{4}}\int_{B_{r}}u^{2}\,dV,$
which via a mean value inequality gives the result.
The previous arguments suggest the possibility of giving an inductive proof of Theorem 2. However, the proof of (1) required the use of Bochner’s formula, and it did not seem to easily translate to an argument in the case of polyharmonic functions (beyond biharmonic ones). Thus, to bypass this stumbling block, we had to give a direct proof of (2) for any polyharmonic function, a genuinely new estimate. This is the most delicate part of our proof, and it nicely sets up the induction argument that seems to be the natural way of proceeding. Thus, although our proof uses the biharmonic case as its basis, it required a new delicate estimate, we believe interesting on its own, for its success.
This result, altogether with Yau’s result and our recent work on biharmonic functions, completes on a large class of manifolds of nonnegative Ricci curvature a natural hierarchy of Yau-Liouville type theorems for the Laplacian and its iterations, and opens up the question of whether a similar phenomenon also holds for elliptic operators of higher order, as it does in $\mathbb{R}^{n}$.
$1.$
The interplay between the behavior of polyharmonic functions and curvature is largely underexplored, and it might reveal some interesting and surprising results. As it is shown in this paper, if $\mbox{Ric}_{g}\geq 0$, the situation is similar to that of the Euclidean spaces. On the other hand, in the case of negative curvature, the authors in [5] have shown that there is a curvature regime where a bounded biharmonic function must be harmonic (and there are bounded nonconstant harmonic functions). So a natural question arises: what happens in the case of $k$-polyharmonic functions?
1.2.
In a previous version of this manuscript, we made the following assumptions
  • (H1)
    the spheres around $p$ are mean-convex with respect to the radial vector field, or
  • (H2)
    The square of the distance function from $p$, that we denote by $d$, is convex.
Any of these assumptions implies a lower bound on the Laplacian of the distance function from the pole, which is needed to construct appropriate cutoff functions. However, and a bit serendipitously, we discovered that the hypothesis $\mbox{Ric}_{g}\geq 0$ plus the existence of a pole implies (H1), so it becomes a redundant assumption (see [6]).
1.3. Organization of the paper
The material of this paper is organized as follows. In Section 2, we collect some facts about cutoff functions and the hole-filling technique needed in the proof of the main $L^{2}$-estimates needed in the proof of our main result. These $L^{2}$ estimates are presented and proved in Section 3.1. Finally, Theorem 2 is proved in Section 3.2.
2. Some Technical Lemmas
In this section, we collect some technical lemmas that will be used in the proof of our main result.
The following lemma is proved in [6].
Lemma 1.
Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$. Set $B_{\rho}:=B_{\rho}(p)$, assume $\mbox{Ric}_{g}\geq 0$ on $B_{R}$, and let $0<r<R$. Then there exists $\varphi\in C_{c}^{\infty}(B_{r})$ such that $0\leq\varphi\leq 1$, $\varphi=1$ on $B_{s}$, and for $\dfrac{1}{8}r<s<r<R$,
$\left|\nabla\varphi\right|^{2}\leq\frac{C}{\left(r-s\right)^{2}}\varphi\quad\text{ and }\quad\frac{\left|\Delta\left(\varphi^{4}\right)\right|^{2}}{\varphi^{4}}\leq\frac{C}{\left(r-s\right)^{4}}.$
In particular, we may assume that the weaker bounds
$|\Delta\varphi|\leq\frac{C}{(r-s)^{2}}\quad\mbox{and}\quad\frac{|\Delta(\varphi^{2})|}{\varphi}\leq\frac{C}{(r-s)^{2}}$
hold and will be used below.
The following lemma and its proof can also be found in [6].
Lemma 2 (Hole-filling iteration).
Let $F,G:[0,R]\to[0,\infty)$ be functions, $\alpha>0$ and assume that there exists $\theta\in[0,1)$ such that for all $0<s<r\leq R$ one has
(3) $F(s)\;\leq\;\theta\,F(r)\;+\;\frac{1}{(r-s)^{\alpha}}\,G(r).$
Suppose further that $G$ is bounded above by $M$. Then for any $\lambda\in(\theta^{1/\alpha},1)$,
(4) $F(s)\;\leq\;\frac{(1-\lambda)^{-\alpha}}{1-\theta\lambda^{-\alpha}}\,\frac{M}{(r-s)^{\alpha}},\qquad 0<s<r\leq R.$
The following basic estimate is Theorem 6.2 in Chapter II from [17].
Lemma 3 (Mean value inequality for $u^{2}$).
Let $(M^{n},g)$ be a complete Riemannian manifold and fix $x_{0}\in M$. Assume a lower Ricci bound $\mbox{Ric}_{g}\geq-(n-1)Kg$ on $B_{2a}:=B_{2a}(x_{0})$ for some $K\geq 0$. If $u\geq 0$ is (weakly) subharmonic on $B_{2a}$ (i.e. $\Delta u\geq 0$ in $B_{2a}$), then
(5) $\sup_{B_{a}}u^{2}\;\leq\;C(n)\,\exp\!\big(C(n)\sqrt{K}\,R\big)\,\frac{1}{\mathrm{Vol}(B_{2a})}\int_{B_{2a}}u^{2}.$
In particular, when $K=0$ (nonnegative Ricci curvature on $B_{2a}$),
(6) $\sup_{B_{a}}u^{2}\;\leq\;\frac{C(n)}{\mathrm{Vol}(B_{2a})}\int_{B_{2a}}u^{2}.$
2.1. An estimate for the Laplacian of biharmonic functions
The following estimate is the basis of our inductive argument in order to prove (2). As discussed in the appendix, it is a consequence of the arguments given in [6].
Lemma 4.
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with $\mbox{Ric}_{g}\geq 0$. Let $a>0$ and let $u\in C^{\infty}(B_{a})$ be Biharmonic, i.e. $\Delta^{2}u=0$ on $B_{a}$. Then, there is a constant $C>0$ such that for every $s,r\in(0,a)$, with $0<\frac{1}{8}r<s<r<a$
(7) $\int_{B_{s}}|\Delta u|^{2}dV\leq\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV.$
3. Main estimates and proof of Theorem 2
3.1. $L^{2}$ estimates
The following energy estimate is valid for any sufficiently smooth function.
Lemma 5.
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with $\mathrm{Ric}_{g}\geq 0$. Let $r,s,a>0$ be such that $s<r<a$, and let $u\in C^{\infty}(B_{a})$, and let $\varphi$ be a cutoff function as in Lemma 1. Then, there exists a constant $C=C(n)>0$ such that
(8) $\int_{B_{r}}|\nabla\Delta u|^{2}\varphi^{2}dV\leq\frac{(r-s)^{2}}{C}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{2}}\int_{B_{r}}|\Delta u|^{2}dV.$
Proof.
Fix $0<s<r<a$. Let $\varphi\in C_{c}^{\infty}(B_{r}(p))$ as in Lemma 1, that is
$0\leq\varphi\leq 1,\qquad\varphi\equiv 1\;\text{ on }B_{s}(p),\qquad\mathrm{supp}\,\varphi\subset B_{r}(p).$
with
(9) $|\nabla\varphi|\leq\frac{C}{r-s},\qquad|\Delta\varphi|\leq\frac{C}{(r-s)^{2}}.$
We shall use the notation $B_{\rho}=B_{\rho}\left(p\right)$.
Let $\Delta u=f$, by integration by parts on $B_{r}$,
(10) $\int_{B_{r}}\varphi^{2}\,|\nabla f|^{2}\,dV$ $=-\int_{B_{r}}\varphi^{2}\,f\,\Delta f\,dV-2\int_{B_{r}}\varphi\,f\,\langle\nabla\varphi,\nabla f\rangle\,dV.$
Applying Young’s inequality to the last term, for any $\varepsilon>0$,
$\left|2\int_{B_{r}}\varphi\,f\,\langle\nabla\varphi,\nabla f\rangle\,dV\right|\leq\varepsilon\int_{B_{r}}\varphi^{2}|\nabla f|^{2}\,dV+\frac{C}{\varepsilon(r-s)^{2}}\int_{B_{r}}f^{2}\,dV,$
where we used $|\nabla\varphi|\leq C/(r-s)$ from (9). Choosing $\varepsilon=\tfrac{1}{2}$ and using (10), and Young’s inequality again with $\delta>0$,
$\int_{B_{r}}\varphi^{2}|\nabla f|^{2}\,dV$ $\leq 2\int_{B_{r}}\varphi^{2}|f\,\Delta f|\,dV+\frac{C}{(r-s)^{2}}\int_{B_{r}}f^{2}\,dV,$
$\leq\delta\int_{B_{r}}f^{2}\varphi^{2}dV+\frac{1}{\delta}\int_{B_{r}}|\Delta f|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{2}}\int_{B_{r}}f^{2}\,dV,$
or
$\int_{B_{r}}\varphi^{2}|\nabla\Delta u|^{2}dV\leq\delta\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV+\frac{1}{\delta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{2}}\int_{B_{r}}|\Delta u|^{2}dV.$
Taking $\delta=\frac{C}{(r-s)^{2}}$ we obtain,
$\int_{B_{r}}\varphi^{2}|\nabla\Delta u|^{2}dV\leq\frac{(r-s)^{2}}{C}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{2}}\int_{B_{r}}|\Delta u|^{2}dV.$
The following Lemma allows to set up the induction argument in the proof of our main result. The proof is by induction, the base case being the biharmonic case (Lemma 4).
Lemma 6.
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with $\mathrm{Ric}_{g}\geq 0$. Fix $a>0$ and let $u\in C^{\infty}(B_{r})$ be k-polyharmonic, i.e. $\Delta^{k}u=0$ on $B_{r}$, for $k\geq 2$. Then, there exists a constant $C>0$ such that for every $s,r$, with $0<\frac{r}{8}<s<r$,
(11) $\int_{B_{s}}|\Delta u|^{2}dV\leq\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV.$
Proof.
We proceed by induction. For $k=2$, the estimate is just 4. Assume that the estimate holds for (k1)(k-1)-polyharmonic functions, with $k\geq 3$.
Fix $0<s<r$. By Lemma 1 we can choose $\varphi\in C_{c}^{\infty}(B_{r}(p))$ such that
$0\leq\varphi\leq 1,\qquad\varphi\equiv 1\;\text{ on }B_{s},\qquad\mathrm{supp}\,\varphi\subset B_{r}.$
with
(12) $|\nabla\varphi|^{2}\leq\frac{C}{(r-s)^{2}}\varphi,\qquad|\Delta\varphi|\leq\frac{C}{(r-s)^{2}},$
and hence
(13) $\frac{|\Delta(\varphi^{2})|}{\varphi}\leq\frac{2\varphi|\Delta\varphi|}{\varphi}+\frac{2|\nabla\varphi|^{2}}{\varphi}\leq\frac{C}{(r-s)^{2}}.$
By Young’s inequality, for $\eta>0$, $\epsilon>0$ and $\delta>0$, we have
$\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV$ $=\int_{B_{r}}u\Delta\left(\varphi^{2}\Delta u\right)dV,$
$=\int_{B_{r}}u\Delta^{2}u\varphi^{2}dV+\int_{B_{r}}u\Delta u\Delta\varphi^{2}dV+4\int_{B_{r}}u\varphi\langle\nabla\varphi,\nabla\Delta u\rangle dV$
$\leq\frac{\eta}{2}\int_{B_{r}}u^{2}\varphi^{2}dV+\frac{1}{2\eta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\int_{B_{r}}\frac{\epsilon}{2}u^{2}dV$
$\quad+\int_{B_{r}}|\Delta u|^{2}\frac{|\Delta\varphi^{2}|^{2}}{2\epsilon}dV+4\int_{B_{r}}\frac{1}{2\delta}u^{2}\varphi^{2}dV+4\int_{B_{r}}\frac{\delta}{2}|\nabla\Delta u|^{2}|\nabla\varphi|^{2}dV.$
Taking $\epsilon=\dfrac{2C}{(r-s)^{4}}$ yields
$\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV$ $\leq\frac{\eta}{2}\int_{B_{r}}u^{2}\varphi^{2}dV+\frac{1}{2\eta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$\quad+\frac{(r-s)^{4}}{4C}\int_{B_{r}}|\Delta u|^{2}|\Delta\varphi^{2}|^{2}dV+\int_{B_{r}}\frac{2}{\delta}u^{2}\varphi^{2}dV$
$\quad+\int_{B_{r}}2\delta|\nabla\Delta u|^{2}|\nabla\varphi|^{2}dV.$
Applying the bounds on the cutoff function given in (12) and (13)
$\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV$ $\leq\frac{\eta}{2}\int_{B_{r}}u^{2}dV+\frac{1}{2\eta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$\quad+\frac{(r-s)^{4}}{4C}\frac{C}{(r-s)^{4}}\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV+\int_{B_{r}}\frac{2}{\delta}u^{2}\varphi^{2}dV$
$\quad+\frac{C}{(r-s)^{2}}\int_{B_{r}}2\delta|\nabla\Delta u|^{2}dV.$
Taking $\delta=\dfrac{(r-s)^{4}}{32C^{2}}\varphi^{2}$,
$\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV$ $\leq\frac{\eta}{2}\int_{B_{r}}u^{2}dV+\frac{1}{2\eta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$\quad+\frac{1}{4}\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV+\frac{64C^{2}}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$\quad+\frac{1}{16}\frac{(r-s)^{2}}{C}\int_{B_{r}}|\nabla\Delta u|^{2}\varphi^{2}dV.$
By Lemma 5 we have that
$\int_{B_{r}}|\Delta u|^{2}\varphi^{2}dV$ $\leq\frac{\eta}{2}\int_{B_{r}}u^{2}dV+\frac{1}{2\eta}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV+\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$\quad+\frac{1}{4}\int_{B_{r}}|\Delta u|^{2}dV+\frac{64C^{2}}{(r-s)^{4}}\int_{B_{r}}u^{2}dV$
$+\frac{1}{16}\frac{(r-s)^{4}}{C^{2}}\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV\quad+\frac{1}{16}\int_{B_{r}}|\Delta u|^{2}dV.$
As $u$ is $k$-polyharmonic then $\Delta u$ is (k1)(k-1)-polyharmonic and thus, by the induction hypothesis,
(14) $\int_{B_{r}}|\Delta^{2}u|^{2}\varphi^{2}dV\leq\frac{C^{2}}{(r-s)^{4}}\int_{B_{r}}|\Delta u|^{2}dV.$
Therefore, using (14) in the previous estimate,
$\int_{B_{s}}|\Delta u|^{2}dV$ $\leq\left[\frac{\eta}{2}+\frac{C}{(r-s)^{4}}+\frac{64C^{2}}{(r-s)^{4}}\right]\int_{B_{r}}u^{2}dV$
$\quad+\left[\frac{1}{2\eta}\frac{C^{2}}{(r-s)^{4}}+\frac{1}{4}+\frac{1}{16}+\frac{1}{16}\right]\int_{B_{r}}|\Delta u|^{2}dV.$
Taking $\eta=\dfrac{4C^{2}}{(r-s)^{4}}$ gives
(15) $\int_{B_{s}}|\Delta u|^{2}dV\leq\frac{C_{n}}{(r-s)^{4}}\int_{B_{r}}u^{2}dV+\frac{1}{2}\int_{B_{r}}|\Delta u|^{2}dV.$
Applying the hole-filling lemma (Lemma 2) to (15) with $\alpha=4$, we have that for $0<\rho\leq r$,
$\int_{B_{s}}|\Delta u|^{2}dV\leq\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}dV.$
We have the following important consequence which follows by induction from the previous lemma.
Corollary 7.
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with $\mathrm{Ric}_{g}\geq 0$. Fix $a>0$ and let $u\in C^{\infty}(B_{r})$ be k-polyharmonic, i.e. $\Delta^{k}u=0$ on $B_{r}$, for $k\geq 2$. Then, there is a constant $C>0$ such that for every $r>0$ it holds that
$\int_{B_{r}}|\Delta^{k-1}u|^{2}dV\leq\frac{2^{4\left(k-2\right)}C}{r^{4(k-1)}}\int_{B_{2^{k-1}r}}u^{2}dV.$
3.2. Proof of Main Theorem 2
We are ready to prove our main result. Let $u$ be a $k$-polyharmonic function. By Corollary 7 we can estimate
$\int_{B_{r}}|\Delta^{k-1}u|^{2}\,dV\leq\frac{C^{\prime}}{r^{4(k-1)}}\int_{B_{2^{k-1}r}}u^{2}\,dV.$
By the volume doubling property (which holds if $\mbox{Ric}\geq 0$), we have
$\frac{1}{\mbox{Vol}\left(B_{r}\right)}\int_{B_{r}}|\Delta^{k-1}u|^{2}\,dV$ $\leq$ $\frac{C^{\prime}}{r^{4(k-1)}}\frac{1}{\mbox{Vol}\left(B_{2^{k-1}r}\right)}\int_{B_{2^{k-1}r}}u^{2}\,dV,$
and hence, since $\Delta^{k-1}u$ is harmonic, then $|\Delta^{k-1}u|$ is subharmonic, by the mean value inequality (Lemma 3),
$\sup_{x\in B_{r/2}}\left|\Delta^{k-1}u\right|^{2}$ $\leq$ $\frac{C^{\prime\prime}}{r^{4(k-1)}}\frac{1}{\mbox{Vol}\left(B_{2^{k-1}r}\right)}\int_{B_{2^{k-1}r}}u^{2}\,dV$
$\leq$ $\frac{C^{\prime\prime}}{r^{4(k-1)}}\sup_{x\in B_{r}\left(p\right)}\left|u\left(x\right)\right|^{2}=\frac{1}{r^{4(k-1)}}o\left(r^{4(k-1)}\right)$
and hence if $u=o(r^{2(k-1)})$, letting $r\rightarrow\infty$, it follows that
$\Delta^{k-1}u\equiv 0,$
i.e., $u$ is (k1)(k-1)-polyharmonic. For the last statement, let $u$ be a polyharmonic function of sublinear growth, i.e.
$\sup_{B_{r}}|u|=o(r)\quad\text{as }r\to\infty.$
For every $k\geq 2$ we have
$\frac{\sup_{B_{r}}|u|}{r^{2(k-1)}}=o\big(r^{1-2(k-1)}\big)\;\longrightarrow\;0.$
In particular $u=o\big(r^{2(k-1)}\big)$ for each $k\geq 2$. Assume that $u$ is kk–polyharmonic for some $k\geq 2$. By the first part of the theorem, $u$ is then (k1)(k-1)–polyharmonic. Iterating this argument for $k,k-1,\dots,2$, we conclude that $u$ is harmonic and it is sublinear growth.
Finally, by Yau’s Liouville theorem for harmonic functions on complete manifolds with $\mathrm{Ric}_{g}\geq 0$, any harmonic function with sublinear growth must be constant. Hence $u$ is constant on $M$, which is what we wanted to prove.
Appendix A The biharmonic case
As stated in the introduction, inequality (2) does not appear in [6]. However, the estimate can be inferred from the proofs of Lemmas 5 and 6 in [6] as follows. From the proof of the local energy estimate for the Laplacian, Lemma 5, for a cutoff function $\varphi\in C_{c}^{\infty}(B_{2r})$ satisfying for $0<\frac{1}{8}r<s<r$
$\varphi\equiv 1\text{ on }B_{s},\quad 0\leq\varphi\leq 1,\quad|\nabla\varphi|\leq\frac{C_{0}}{r-s}\varphi,\quad|\Delta\varphi|\leq\frac{C_{0}}{(r-s)^{2}},$
we have
Lemma 8.
Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with a pole $p$ with $\mathrm{Ric}_{g}\geq 0$. Let $B_{\rho}\colon=B_{\rho}\left(p\right)$. Fix $a>0$ and let $u\in C^{\infty}(B_{a})$ be biharmonic, i.e. $\Delta^{2}u=0$ on $B_{2a}$. Then there exists a constant $C=C(n)>0$ such that, for every pair of radii $0<\dfrac{1}{8}r<s<r\leq a$,
$\int_{B_{s}}|\Delta u|^{2}\,dV\leq\frac{C}{(r-s)^{2}}\int_{B_{r}}|\nabla u|^{2}\varphi^{2}\,dV\;+\;\frac{1}{2}\int_{B_{r}}|\Delta u|^{2}\,dV.$
From this lemma, using a hole-filling technique follows that for every $0<\dfrac{1}{8}r<s<r<\infty$
$\int_{B_{s}}|\Delta u|^{2}\,dV\leq\frac{C}{(r-s)^{2}}\int_{B_{r}}|\nabla u|^{2}\varphi^{2}\,dV.$
From the proof of Lemma 6, the Cacciopoli inequality for a biharmonic function, for a cutoff $\varphi\in C_{c}^{\infty}(B_{2r})$ such that $0\leq\varphi\leq 1$, $\varphi\equiv 1$ on $B_{s}$, and for $\dfrac{1}{8}r<s<r$, it also satisfies
$\left|\nabla\varphi\right|^{2}\leq\frac{C}{\left(r-s\right)^{2}}\varphi\quad\text{ and }\quad\frac{\left|\Delta\left(\varphi^{4}\right)\right|^{2}}{\varphi^{4}}\leq\frac{C}{\left(r-s\right)^{4}},$
we have
Lemma 9.
Let $(M,g)$ be a smooth Riemannian manifold with a pole $p\in M$, such that $\mbox{Ric}\geq 0$. Let $0<\dfrac{1}{8}r<s<r<\infty$ and let $B_{\rho}:=B_{\rho}(p)$. If $u\in C^{\infty}(B_{2r})$ is biharmonic, i.e. $\Delta^{2}u=0$ in $B_{2r}$, then there exists a constant $C=C(n)>0$ such that
$\int_{B_{r}}|\nabla u|^{2}\varphi^{2}\,dV\;\leq\;\frac{C}{(r-s)^{2}}\int_{B_{r}}u^{2}\,dV.$
Putting all these together, (2) follows readily.

### v2

@@
 Abstract.
-In this paper, we establish a complete Liouville–type hierarchy for polyharmonic functions on a broad class of Riemannian manifolds with nonnegative Ricci curvature.
-Extending Yau’s classical result for harmonic functions and our recent biharmonic Liouville theorem, we prove that on any complete manifold of nonnegative Ricci curvature with a pole, every kk–polyharmonic function of growth $o(r^{2(k-1)})$ must in fact be (k1)(k-1)–polyharmonic.
+In this paper, we establish a complete Liouville–type hierarchy for polyharmonic functions on Riemannian manifolds with nonnegative Ricci curvature.
+Extending Yau’s classical result for harmonic functions and our recent biharmonic Liouville theorem, we prove that on any complete manifold of nonnegative Ricci curvature, every kk–polyharmonic function of growth $o(r^{2(k-1)})$ must in fact be (k1)(k-1)–polyharmonic.
 Iterating this procedure yields the result that all polyharmonic functions of sublinear growth are constant.
@@
 1. Introduction
-Harmonic functions have been a central object of study in Geometric Analysis.
-Among the most celebrated results, and a foundational one, is Yau’s Liouville theorem [18], from 1975, which asserts that on a complete Riemannian manifold of nonnegative Ricci curvature, any bounded harmonic function must be constant.
-Since Yau, the number of interesting and important results about harmonic functions and spaces of harmonic functions have been constantly increasing.
-Among them, just to name a few, there are the beautiful works of Li and Tam [13], Colding and Minicozzi [7, 8] on the structure of the spaces of harmonic functions of polynomial growth, Ni and Tam [15] on pluriharmonic functions, and more recently, the work of Liu for holomorphic functions in [14].
-However, what might seem a natural question has received little attention in the literature.
-Although the intrinsic interest the study of higher order elliptic operators may have (for instance, the Paneitz operator in conformal, or the study of polyharmonic flows for curves) there are very few results regarding their Liouville-type properties.
-As a few interesting exceptions we refer to the works of Branding [2, 3], where the author considers biharmonic maps, and proves Liouville-type theorems under the smallness of certain Sobolev norms.
-Recently the authors proved the following theorem (see [6]).
+Harmonic functions have long been a central object of study in geometric analysis.
+Among the most celebrated and foundational results in the field is Yau’s Liouville theorem [20], which asserts that on a complete Riemannian manifold with nonnegative Ricci curvature, any bounded harmonic function must be constant.
+This theorem not only captures a fundamental rigidity phenomenon, but also inaugurated a far-reaching program aimed at understanding how analytic growth conditions reflect the large-scale geometry of the underlying manifold.
+Since Yau’s work, the study of harmonic functions and spaces of harmonic functions has developed extensively.
+Notable contributions include the work of Li and Tam [14], as well as Colding and Minicozzi [8, 9], who investigated the structure and finite dimensional properties of spaces of harmonic functions of polynomial growth.
+Further important developments include results of Ni and Tam [16] on pluriharmonic functions, and more recently Liu’s work on holomorphic functions via three–circle theorems [15].
+Collectively, these results demonstrate that under nonnegative Ricci curvature, harmonic functions exhibit rigidity properties closely resembling those of the Euclidean setting.
+By contrast, Liouville-type phenomena for higher-order elliptic equations have received comparatively little attention.
+This gap is striking given the intrinsic analytic and geometric relevance of higher-order operators, such as the Paneitz operator in conformal geometry or the operators arising in the study of polyharmonic flows for curves.
+Aside from a few notable exceptions, there are very few results addressing Liouville-type properties in this broader context.
+Among these, we mention the works of Branding [3, 4], where Liouville-type theorems are obtained for biharmonic maps under smallness assumptions on certain Sobolev norms.
+Recently, the authors proved the following result (see [7], as well as the related work [19]).
 Theorem 1.
 Let $(M,g)$ be a complete Riemannian manifold with a pole and such that $\mbox{Ric}_{g}\geq 0$.
-Then any biharmonic function $u\in C^{4}\left(M\right)$ of subquadratic growth is harmonic.
-Therefore, any biharmonic function of sublinear growth must be constant.
-The purpose of this paper is to extend the previous result to $k$-polyharmonic functions.
-Recall that a $k$-polyharmonic function on a Riemannian manifold $\left(M,g\right)$ is a function $u:M\longrightarrow\mathbb{R}$ is a function $u\in C^{2k}\left(M\right)$ such that
+Then any biharmonic function $u\in C^{4}(M)$ of subquadratic growth is harmonic.
+In particular, any biharmonic function of sublinear growth must be constant.
+The purpose of the present paper is to extend Theorem 1 to the general polyharmonic setting.
+Recall that a $k$-polyharmonic function on a Riemannian manifold $(M,g)$ is a function $u\in C^{2k}(M)$ satisfying
 $\Delta^{k}u=0,$
-where $\Delta$ is the Laplacian.
-In this paper, we shall consider manifolds with a pole, that is manifolds $M$ for which there is a point $p\in M$ such that the exponential map
-$\mbox{exp}_{p}:\,T_{p}M\longrightarrow M$
-is a diffeomorphism We are ready to state our main result.
+where $\Delta$ denotes the Laplacian.
+Our main result is the following.
 Theorem 2.
-Let $(M,g)$ be a complete Riemannian manifold with a pole and such that $\mbox{Ric}_{g}\geq 0$. Then any $k$-polyharmonic function $u\in C^{2k}\left(M\right)$ with
+Let $(M,g)$ be a complete Riemannian manifold with $\mbox{Ric}_{g}\geq 0$. Then any $k$-polyharmonic function $u\in C^{2k}(M)$ satisfying
 $u=o(r^{2(k-1)})\quad\mbox{as}\quad r\rightarrow\infty,$
-where $r=d\left(p,x\right)$, is (k1)(k-1)-polyharmonic. Therefore, any polyharmonic function of sublinear growth must be constant.
-As the reader may know, Theorem 2 has been known in the case of $\mathbb{R}^{n}$ with the Euclidean metric for quite some time: the critical growth separating genuinely $k$-polyharmonic behavior from lower-order behavior is $r^{2(k-1)}$.
-This shows that our theorem is sharp and is a natural extension of what is known in the Euclidean setting.
-We now give an account of the subtleties in the proof of Theorem 2.
-First, we review the proof of Theorem 1.
-Our proof in [6] used an energy estimate for the $L^{2}$-norm of $\Delta u$ and a Cacciopoli-type inequality
-(1) $\int_{B_{s}}\left|\nabla u\right|^{2}\,dV\leq\frac{C}{\left(r-s\right)^{2}}\int_{B_{r}}u^{2}\,dV$
-valid for both harmonic and biharmonic functions. All this was used to prove
-(2) $\int_{B_{s}}\left|\Delta u\right|^{2}\,dV\leq\frac{C}{\left(r-s\right)^{4}}\int_{B_{r}}u^{2}\,dV,$
-which via a mean value inequality gives the result.
-The previous arguments suggest the possibility of giving an inductive proof of Theorem 2.
-However, the proof of (1) required the use of Bochner’s formula, and it did not seem to easily translate to an argument in the case of polyharmonic functions (beyond biharmonic ones).
-Thus, to bypass this stumbling block, we had to give a direct proof of (2) for any polyharmonic function, a genuinely new estimate.
-This is the most delicate part of our proof, and it nicely sets up the induction argument that seems to be the natural way of proceeding.
-Thus, although our proof uses the biharmonic case as its basis, it required a new delicate estimate, we believe interesting on its own, for its success.
-This result, altogether with Yau’s result and our recent work on biharmonic functions, completes on a large class of manifolds of nonnegative Ricci curvature a natural hierarchy of Yau-Liouville type theorems for the Laplacian and its iterations, and opens up the question of whether a similar phenomenon also holds for elliptic operators of higher order, as it does in $\mathbb{R}^{n}$.
+where $r=d(p,x)$, is (k1)(k-1)-polyharmonic. In particular, any polyharmonic function of sublinear growth must be constant.
+As is well known, Theorem 2 has long been established in the Euclidean setting.
+In $\mathbb{R}^{n}$, the critical growth rate that separates genuinely $k$-polyharmonic behavior from lower-order behavior is precisely $r^{2(k-1)}$, as follows from Almansi’s decomposition.
+This shows that Theorem 2 is sharp and constitutes a natural geometric extension of the classical Euclidean theory.
+We now briefly describe the main analytic difficulty underlying the proof of Theorem 2.
+We begin by recalling the strategy used in the proof of Theorem 1.
+In [7], the argument relied on an energy estimate for the $L^{2}$-norm of $\Delta u$ together with a Caccioppoli-type inequality,
+(1) $\int_{B_{s}}|\nabla u|^{2}\,dV\leq\frac{C}{(r-s)^{2}}\int_{B_{r}}u^{2}\,dV,$
+valid for both harmonic and biharmonic functions. This estimate was then used to derive
+(2) $\int_{B_{s}}|\Delta u|^{2}\,dV\leq\frac{C}{(r-s)^{4}}\int_{B_{r}}u^{2}\,dV,$
+which, combined with a mean value inequality, yields the desired rigidity.
+At first glance, these arguments suggest the possibility of an inductive approach to Theorem 2.
+However, the proof of (1) is fundamentally based on the Bochner formula, and this approach does not extend in any straightforward way to polyharmonic functions beyond the biharmonic case.
+To overcome this obstacle, we establish a direct analogue of (2) for arbitrary polyharmonic functions.
+This new estimate is the most delicate part of the argument and forms the analytic backbone of the inductive procedure.
+Although the biharmonic case serves as the base step, the proof requires a genuinely new estimate that may be of independent interest.
+Taken together with Yau’s theorem and our previous work on biharmonic functions, the results of this paper complete, on manifolds with nonnegative Ricci curvature, a natural hierarchy of Yau–Liouville type theorems for the Laplacian and its iterates.
+They also raise the question of whether similar rigidity phenomena persist for more general higher-order elliptic operators, as they do in the Euclidean setting.
+Beyond its analytic content, Theorem 2 may be interpreted as a rigidity principle for complete manifolds with nonnegative Ricci curvature.
+It shows that, at the level of critical growth, such manifolds are analytically indistinguishable from Euclidean space with respect to iterates of the Laplacian.
+In particular, the behavior of polyharmonic function in nonnegative Ricci curvature seems to mirror Almansi’s decomposition in $\mathbb{R}^{n}$.
 $1.$
-The interplay between the behavior of polyharmonic functions and curvature is largely underexplored, and it might reveal some interesting and surprising results.
-As it is shown in this paper, if $\mbox{Ric}_{g}\geq 0$, the situation is similar to that of the Euclidean spaces.
-On the other hand, in the case of negative curvature, the authors in [5] have shown that there is a curvature regime where a bounded biharmonic function must be harmonic (and there are bounded nonconstant harmonic functions).
-So a natural question arises: what happens in the case of $k$-polyharmonic functions?
+The interplay between the behavior of polyharmonic functions and curvature remains largely unexplored and may reveal further rigidity phenomena.
+As shown in this paper, when $\mbox{Ric}_{g}\geq 0$, the qualitative behavior closely parallels that of Euclidean space.
+In contrast, in negatively curved settings the situation is markedly different.
+In particular, the authors showed in [6] that there exist curvature regimes in which every bounded biharmonic function must be harmonic, while nonconstant bounded harmonic functions still exist.
+This naturally leads to the question of how $k$-polyharmonic functions behave in such settings.
 1.2.
-In a previous version of this manuscript, we made the following assumptions
+In a previous version of this manuscript, we imposed one of the following additional assumptions:
   • (H1)
-    the spheres around $p$ are mean-convex with respect to the radial vector field, or
+    the geodesic spheres centered at $p$ are mean-convex with respect to the radial vector field, or
   • (H2)
-    The square of the distance function from $p$, that we denote by $d$, is convex.
-Any of these assumptions implies a lower bound on the Laplacian of the distance function from the pole, which is needed to construct appropriate cutoff functions.
-However, and a bit serendipitously, we discovered that the hypothesis $\mbox{Ric}_{g}\geq 0$ plus the existence of a pole implies (H1), so it becomes a redundant assumption (see [6]).
+    the square of the distance function from $p$, denoted by $d$, is convex.
+Either assumption yields a lower bound for the Laplacian of the distance function, which is needed to construct suitable cutoff functions.
+However, we later observed that the condition $\mbox{Ric}_{g}\geq 0$, together with the existence of a pole, already implies (H1), rendering it redundant (see [7]).
+Subsequently, inspired by the recent work [19] and with the generous assistance of Professor Wang, we adopted the cutoff technology developed in [2], which requires only the assumption $\mbox{Ric}_{g}\geq 0$.
+We are grateful to Professors Wang and Zhu for their help.
 1.3. Organization of the paper
@@
 In this section, we collect some technical lemmas that will be used in the proof of our main result.
-The following lemma is proved in [6].
+The following pair of lemmas are proved in [7] under the hypothesis that the manifold has a pole.
+However, using the cutoff technology from [2], we can give a proof of its validity without this extra assumption.
+Lemma 1 (Cutoff with explicit bounds).
+Let $(M^{n},g)$ be a complete Riemannian manifold with $\mbox{Ric}_{g}\geq 0$ Set $B_{\rho}:=B_{\rho}(p)$.
+Let $r$ be such that and for $0<\frac{1}{8}R\leq r\leq R$.
+Then there exists $\chi\in C_{c}^{\infty}(B_{R})$ such that
+$0\leq\chi\leq 1,\qquad\chi\equiv 1\ \text{on }B_{r},\qquad\mbox{supp}\left(\chi\right)\subset B_{R},$
+and
+$\left|\nabla\chi\right|\leq\frac{C(n)}{R-r},\quad|\Delta\chi|\leq\frac{C(n)}{(R-r)^{2}}\quad\text{on}\quad B_{R}.$
+Proof.
+Take $\chi=\phi$, where $\phi$ is as in Corollary 2.3 in [2].
+Notice that given $r$, we can write $R=\gamma r$ with $\gamma>1$, and in this case, we obtain the following estimates for the gradient and Laplacian of the cutoff function.
+$\left|\nabla\chi\right|\leq\frac{C_{1}}{R}=\frac{\left(\gamma-1\right)C_{1}}{\left(\gamma-1\right)R}\leq\frac{7C_{1}}{R-r},$
+where we used that $r\geq\frac{1}{8}R$.
+Regarding the Laplacian, as we are in the case $\alpha=2$, we have again that
+$\left|\Delta\chi\right|\leq\frac{C_{2}}{R^{2}}\leq\frac{\left(\gamma-1\right)^{2}C_{2}}{\left(\gamma-1\right)^{2}R^{2}}\leq\frac{49C_{2}}{\left(R-r\right)^{2}},$
+and the lemma is proved. ∎
+As an immediate consequence we have (see [7]).
 Lemma 2.
-Let $(M^{n},g)$ be a complete Riemannian manifold with a pole $p\in M$.
+Let $(M^{n},g)$ be a complete Riemannian manifold.
 Set $B_{\rho}:=B_{\rho}(p)$, assume $\mbox{Ric}_{g}\geq 0$ on $B_{R}$, and let $0<r<R$.
@@
 Proof.
-We proceed by induction. For $k=2$, the estimate is just 4. Assume that the estimate holds for (k1)(k-1)-polyharmonic functions, with $k\geq 3$.
+We proceed by induction. For $k=2$, the estimate is just 5. Assume that the estimate holds for (k1)(k-1)-polyharmonic functions, with $k\geq 3$.
 Fix $0<s<r$. By Lemma 2 we can choose $\varphi\in C_{c}^{\infty}(B_{r}(p))$ such that
@@
 Lemma 9.
-Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with a pole $p$ with $\mathrm{Ric}_{g}\geq 0$.
+Let $(M,g)$ be an nn–dimensional complete Riemannian manifold with $\mathrm{Ric}_{g}\geq 0$.
 Let $B_{\rho}\colon=B_{\rho}\left(p\right)$.
@@
 Lemma 10.
-Let $(M,g)$ be a smooth Riemannian manifold with a pole $p\in M$, such that $\mbox{Ric}\geq 0$.
+Let $(M,g)$ be a smooth Riemannian manifold such that $\mbox{Ric}\geq 0$.
 Let $0<\dfrac{1}{8}r<s<r<\infty$ and let $B_{\rho}:=B_{\rho}(p)$.